OpenStudy (anonymous):

Solve the equation: 3.3n = 12.2n + 21.6 A)1.5 B)2.1 C) -1.5 D) -2.1

OpenStudy (anonymous):

So start with getting rid of the same terms.

OpenStudy (anonymous):

Oh sorrry there is no 3.3n .. only 3.3 lol fail

OpenStudy (anonymous):

oh thats alright lemme change some of my math real quick

OpenStudy (anonymous):

Okie. I got -18.3 = 12.2 n

OpenStudy (anonymous):

good job! Next diving -18.3 by 12.2 and tell me what you get

OpenStudy (anonymous):

I have the answer i just want to know if you can get it.

OpenStudy (anonymous):

I got 1.5 :)

OpenStudy (anonymous):

Close but wrong.

OpenStudy (anonymous):

What happens why you divide a negative by a positive?

OpenStudy (anonymous):

Oh it becomes a Negative!

OpenStudy (anonymous):

The correct answer is -1.5 because you keep the negative. Great job other than that! Always keep an eye out for those negatives.

OpenStudy (anonymous):

so -1.5

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Welcome! feel free to tag me in any of your other questions i will try to help out.

OpenStudy (anonymous):

My next question is solve the equation : 3/4 + 1/2x = 7/8 A) x = 1/2 B) x = 1/4 C) x = 3/4

OpenStudy (anonymous):

Show me how you would solve it.

OpenStudy (anonymous):